OpenStudy (anonymous):

210/252= 5/6? reduce to lowest terms...check work

OpenStudy (anonymous):

yes,correct

OpenStudy (anonymous):

210 / 252 = (2*3*5*7) / (2*2*3*3*7) = 5 / (2*3) = 5/6 Yes, that is correct.